OpenStudy (anonymous):

integrate -9 to 0 (2+sqrt of 81-x^2)dx, without using an antiderivative

myininaya (myininaya):

\[\int\limits_{-9}^{0}(2+\sqrt{81}-x^2) dx ? \]

myininaya (myininaya):

\[\int\limits_{-9}^{0}(2+\sqrt{81-x^2})dx\] if it is this one we can do a circle thingy :)

OpenStudy (aum):

I suspect it is the latter.

myininaya (myininaya):

well for the first little part rectangle second part circle \[\int\limits_{-9}^{0}2dx+\int\limits_{-9}^{0}\sqrt{81-x^2} dx\]

myininaya (myininaya):

@xjasmineg are you there?

OpenStudy (anonymous):

yeah, i'm here!

OpenStudy (anonymous):

and its the second one!

OpenStudy (anonymous):

it says evaluate the exact value using area?!

OpenStudy (anonymous):

@myininaya ?

OpenStudy (aum):

The first definite integral is the area of a rectangle with height 2 units and length 9 units. The second one is the area of a quarter of a circle of radius 9.

OpenStudy (aum):



OpenStudy (aum):